solve the equation for h
A= 1/2 bh

Answer by Alan3354(69443)

solve the equation for h
A= 1/2 bh
==================
If it's A = (1/2)bh:
Multiply by 2
Divide by b
-------------
If it's A = 1/(2bh):
Invert
1/A = 2bh
Divide by 2b
1/(2Ab) = h
===========================
Parentheses are free. Use some.
Or A= 1/2 bh ----> A = bh/2